# Pre-training Large Memory Language Models with Internal and External Knowledge

### v2

@@ 3.4 Machine Unlearning: LmLm Supports Instant Forgetting
Figure 6: Evaluation of Machine Unlearning. We compare LmLm with NPO on the TOFU benchmark. Unlike prior methods, LmLm performs unlearning without any additional training. (a–b) Forget quality vs. utility trade-off. LmLm achieves ideal forgetting ($p$-value $>0.05$) without sacrificing general utility. (c–d) LmLm retains knowledge outside the forget set, unlike other methods that degrade retain-set performance due to parameter entanglement. Y-axis labels use -- or $\uparrow$ to indicate retention or improvement. For metrics like Forget ROUGE, where lower isn’t always better, we mark the retain model’s performance as reference.
One natural benefit of decoupling knowledge from model parameters is that editing and unlearning are achievable through simple operations on the external memory, without compromising the model’s general capabilities. To verify this, we extend our evaluation to a standard machine unlearning benchmark, TOFU [33].
@@
We use LLaMa-3.2-1B-Instruct as the base model and compare against NPO [60], a state-of-the-art unlearning method. For LmLm, we perform unlearning by simply removing entries in the database corresponding to the Forget Set. See Appendix B.2 for implementation details.
@@
Unlearning Results. Figure 6 presents the results on the TOFU benchmark, where the Forget Set is 5% of the data. In Figure 6 (a-b), we show the forget quality and model utility throughout the unlearning process, where ideal performance is defined by effective forgetting without degrading model utility. LmLm achieves precisely this—effective forgetting with no loss in model utility—a direct benefit of decoupling factual knowledge from model parameters. Figure 6 (c–d) shows that LmLm preserves knowledge outside the Forget Set, whereas previous training-based methods tend to forget related information due to parameter entanglement. Importantly, forgetting is performed through simple database operations, without model updates or access to the Retain Set. In contrast, the strongest comparable baseline, NPO, incurs a utility degradation, while other unlearning methods either fail to forget or exhibit catastrophic drops in model utility (See Figure 4, right). While LmLm incurs upfront costs for data annotation and model training, it provides substantial payoff in use cases where knowledge editing, removal, or compliance with data deletion requests are necessary.
4 Discussion
Toward Efficient Scaling via Knowledge Offloading.
We summarize our main findings in Figure 4: (1) during pretraining, LmLm consistently obtains lower perplexity compared to its Standard counterparts. (2) LmLm achieves higher factual precision without compromising NLU performance. (3) LmLm supports instant forgetting while maintaining general model capabilities. In particular, even smaller instances of LmLm match or exceed the performance of much larger off-the-shelf models. These results highlight the potential of LmLm to scale efficiently by offloading knowledge to an external database, thereby maintaining strong factual accuracy with fewer parameters. While our experiments are limited to modest scales due to computational constraints, the observed trends suggest that LmLms offer a promising direction toward parameter-efficient language models that externalize factual storage. Such models may enable real-time, verifiable knowledge updates and open up new possibilities for deploying LmLm in resource-constrained or fact-sensitive environments.
@@ 4 Discussion
Figure 7: Training loss on return value tokens.
As shown in Figure 7, models trained with a standard SFT objective quickly reduce the loss on these tokens, suggesting memorization in the model’s parameters. In contrast, LmLm, trained with the masked loss maintains a high loss throughout training, indicating that these facts are not stored internally. LmLm’s successful application in the machine unlearning benchmark (Sec. 3.4) further supports this finding. Additionally, we observe a notable performance gap in Table 3, where factual precision drops substantially when the external database is disabled—forcing the model to rely on its internal parameters. These findings suggest that editing the database offers a direct way to control what the model knows and forgets.
Beyond Entity-Level Knowledge Offloading.
@@
Model Type FactScore (%) ↑ T-REx EM (%) ↑
@@
By leveraging knowledge triplets, LmLm focuses on separating entity-level factual knowledge from linguistic competency. Current research on probing [44, 21], evaluating [36, 55], and editing [35] internal knowledge in language models similarly concentrates on entity-level atomic facts. However, extending our method beyond entity-level knowledge remains a challenge. Under the LmLm framework, this difficulty arises in determining effective formats for separating knowledge beyond entity-focused triplets. For instance, using a simple QA function call format raises concerns about the potential for hallucinated facts during annotation.
@@
LmLm is a promising step toward separating factual knowledge from language models with many exciting future directions. Our current limitations include: (1) It does not guarantee perfect factuality during generation. Noise in the database and errors from fuzzy matching can introduce inaccuracies. However, such issues are easily traceable and verifiable for LmLm. (2) LmLm introduces additional tokens for lookup queries, which increases training and inference costs. (3) The current implementation focuses on entity-level factual knowledge, which captures only a subset of the broader factual spectrum. A more systematic framework is needed to determine which types of knowledge should be delegated to external databases. (4) Our experiments are limited to small models and datasets due to compute constraints. While sufficient to show core benefits, scaling up may improve performance and support more complex reasoning [17], which remains an open challenge for LmLm.
5 Related Work
Parametric vs Non-Parametric Knowledge.
Language models encode knowledge in two main forms: parametric, learned in model weights during training, and non-parametric, retrieved from external sources [34]. Early work showed that pre-trained models [11, 44] implicitly store factual knowledge. Scaling model size further improves factual recall [6, 24], but recent studies highlight fundamental limitations of parametric storage, including hallucination, staleness, poor attribution, and limited adaptability [42, 22, 26, 2], motivating the shift toward non-parametric approaches.
Retrieval-Augmented LLMs.
Non-parametric methods enhance language models with external sources to improve factuality and controllability. Retrieval-augmented generation (RAG)[29, 23, 52] retrieves relevant passages at inference time, while tool-augmented methods[50, 56] enable symbolic reasoning via external API calls. Retrieval-based pretraining approaches such as REALM[19] and RETRO[5] show that incorporating retrieval during training improves generalization and reduces memorization. More recently, SILO [37] demonstrates that even when excluding certain sources from pretraining, competitive perplexity can be maintained through proper retrieval mechanisms. Semi-parametric language models aim to decouple knowledge storage from generation by attaching non-parametric memories of key-value pairs. A prominent example is kNN-LM [27], which retrieves nearest-neighbor training examples at test time to adjust token probabilities without retraining. Follow-up works extend this idea to various settings [43, 28, 51, 57, 40]. Our approach builds on this line of work by focusing on scalable factual memory and dynamic usage during pretraining. Rather than storing full documents or relying on retrieval pipelines at inference time, our method enables the model to learn when to issue structured lookup calls and how to integrate the results into generation.
Machine Unlearning.
Machine unlearning aims to remove specific knowledge from a trained language model without full retraining [25, 7]. The goal is to forget a targeted subset of information while preserving the model’s general capabilities. Existing approaches fall into several main families. Gradient Ascent (GA) increases the loss on the forget set, while Gradient Difference (GD) jointly preserves performance on a retain set [32]. Preference-based methods, such as IdkDPO [33], NPO [60], and SimNPO [15], encourage the model to respond with neutral or deflective answers [47]. Among these, NPO achieves the strongest results on the TOFU benchmark [33], which evaluates entity-level forgetting and utility retention on a synthetic dataset of author profiles.
Despite strong results on TOFU, existing methods struggle to scale to larger forget sets (e.g., 5% of training data) and often degrade model utility or erase unintended related knowledge. These limitations arise from the entanglement of factual content and general linguistic ability in model parameters. Our approach takes a different route by structurally decoupling factual knowledge from the model. By offloading facts to an external database, forgetting reduces to removing entries—offering a scalable, verifiable, and precise solution without compromising utility. Moreover, LmLm can also be applied to knowledge editing benchmarks [38, 53, 16], where modifying or deleting facts becomes as simple as updating the database.
@@ Acknowledgments
SZ is supported by the Defense Advanced Research Projects Agency (DARPA) under Grant No. D24AP00259-00. CKB is supported by the National Science Foundation (NSF) through the NSF Research Traineeship (NRT) program under Grant No. 2345579. JPZ is supported by a grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) (567916). This work is partially supported by LinkedIn-Cornell Partnership; Open Philantropy; the National Science Foundation NSF under awards OAC-2311521, IIS-2107161, IIS-1724282, and HDR-2118310; NASA under award No. 20-OSTFL20-0053; the Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875); DARPA; arXiv; and the New York Presbyterian Hospital. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or of NASA.
@@ A.1 Knowledge Extraction Details
We construct a high-quality seed dataset by sampling 1,000 passages from SQuAD-v2 [48] and 1,000 passages from Wikipedia. These passages are annotated by GPT-4o with structured factual triples and used to train the Annotator model. The remaining Wikipedia passages are used as the pre-training corpus for LmLm, with no overlap with the validation set.
@@
  • The Corrector is fine-tuned using LoRA (r=32, alpha=16[20] for 2 epochs on 19k GPT-4o-annotated SQuAD-v2 passages, with a learning rate of $2\times 10^{-4}$ and an effective batch size of 32. Sequence packing is enabled to improve training efficiency.
@@ A.3 Formalization of Training and Evaluation Objectives
  • $\mathcal{I}_{\text{train}}$: Return values and <|db_end|> are excluded from training loss objective.
@@
$\mathcal{I}_{\text{train}}$:
@@
[…]
@@
    $\text{PPL}_{\text{static/dynamic}}=\exp\left(-\frac{1}{|\mathcal{I}_{\text{org}}|}\sum_{t\in\mathcal{I}_{\text{org}}}\log p_{\theta}(x_{t}\mid x_{<t})\right).$
@@
    $\text{PPL}_{\text{norm}}=\exp\left(-\frac{1}{|\mathcal{T}_{\text{org}}|}\sum_{t\in\mathcal{I}_{\text{train}}}\log p_{\theta}(x_{t}\mid x_{<t})\right)$
@@
    $\text{NLL}(x)=-\sum_{t\in\mathcal{I}_{\text{train}}}\log p_{\theta}(x_{t}\mid x_{<t}).$
@@ A.4 Database and Retrieval Setting
We build our database by annotating the pre-training data, obtaining 54.6M knowledge triplets consisting of 9.5M unique entities, 8.5M relationships and 16.2M unique values. For retrieval, we employ a fuzzy matching mechanism based on the cosine similarity of sentence embeddings from all-MiniLM-L6-v2333https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. As shown in Table 8, specifically, given a lookup call, we compute its embedding and compare it with the embeddings of stored triplets in our database. If the highest similarity score is below a threshold of 0.6, we return unknown to indicate that no sufficiently similar match was found. Alternatively, we implement a prefix-tree constrained generation, ensuring that lookup calls remain covered by the structured knowledge representations. See detailed discussion in Appendix D.3.

### v3

@@ 4.3 Machine Unlearning: LmLm Supports Instant Forgetting
One natural benefit of decoupling knowledge from model parameters is that editing and unlearning are achievable through simple operations on the external memory, without compromising the model’s general capabilities. To verify this, we extend our evaluation to a standard machine unlearning benchmark, TOFU (Maini et al., 2024).
@@
We use LLaMa-3.2-1B-Instruct as the base model and compare against NPO (Zhang et al., 2024), a state-of-the-art unlearning method. For LmLm, we perform unlearning by simply removing entries in the database corresponding to the Forget Set. See Appendix B.2 for implementation details.
Figure 6: Evaluation of Machine Unlearning. We compare LmLm with NPO on the TOFU benchmark. Unlike prior methods, LmLm performs unlearning without any additional training. (a–b) Forget quality vs. utility trade-off. LmLm achieves ideal forgetting ($p$-value >0.05>0.05) without sacrificing general utility. (c–d) LmLm retains knowledge outside the forget set, unlike other methods that degrade retain-set performance due to parameter entanglement. Y-axis labels use -- or $\uparrow$ to indicate retention or improvement. For metrics like Forget ROUGE, where lower isn’t always better, we mark the retain model’s performance as reference.
@@
Unlearning Results. Figure 6 presents the results on the TOFU benchmark, where the Forget Set is 5% of the data. In Figure 6 (a-b), we show the forget quality and model utility throughout the unlearning process, where ideal performance is defined by effective forgetting without degrading model utility. LmLm achieves precisely this—effective forgetting with no loss in model utility—a direct benefit of decoupling factual knowledge from model parameters. Figure 6 (c–d) shows that LmLm preserves knowledge outside the Forget Set, whereas previous training-based methods tend to forget related information due to parameter entanglement. Importantly, forgetting is performed through simple database operations, without model updates or access to the Retain Set. In contrast, RL-based methods such as NPO, a state-of-the-art baseline on TOFU, incur a utility degradation, while other unlearning methods either fail to forget or exhibit catastrophic drops in model utility (See Figure 4, right). While LmLm incurs upfront costs for data annotation and model training, it provides substantial payoff in use cases where knowledge editing, removal, or compliance with data deletion requests are necessary.
4.4 Externalizing Knowledge Improves Factual Precision
Table 2: Evaluations on factual precision. Subscripts show absolute difference from their respective Standard baselines.
Model Model Type FactScore ↑ T-REx EM ↑ PopQA Acc ↑
OpenAI/GPT2-124M - 14.6 20.1 18.51
GPT2-124M Standard 10.7 41.2 18.51
LmLm 20.6+9.9 54.6+13.4 49.89+31.4
LLaMA2-176M Standard 10.1 46.3 24.59
LmLm 30.6+20.5 54.1+7.8 49.61+25.0
OpenAI/GPT2-355M - 15.2 28.4 19.1
GPT2-355M Standard 14.4 44.9 21.4
LmLm 23.9+9.5 58.7+13.8 52.0+30.6
LLaMA2-382M Standard 14.0 52.0 22.7
LmLm 31.9+17.9 58.1+6.1 50.8+28.1
OpenAI/GPT2-774M - 17.4 35.6 18.5
Pythia-1B - 21.1 47.8 19.5
LLaMA2-7B - 34.0 60.5 29.2
LLaMA3.1-8B - 40.3 67.3 29.4
We evaluate factual precision on long-form biography generation (FactScore) (Min et al., 2023) and short-form factual completion using T-REx (Petroni et al., 2021) and the long-tail subset of PopQA (Asai et al., 2023). FactScore measures the proportion of atomic facts supported by a trusted source, while T-REx and PopQA follow EM and Acc metrics from prior work (Schick et al., 2023; Asai et al., 2023).
As shown in Table 2, LmLm consistently improves factual precision over Standard, its knowledge-dense counterpart. On LLaMA2-382M, it gains $17.9\%$ FactScore, $6.1\%$ T-REx EM, and $28.1\%$ PopQA Acc. Notably, LmLm approaches the performance of much larger models such as Pythia-1B and LLaMA2-7B, despite using far fewer parameters.
How Does LmLm Compare to RAG?
Unlike Retrieval-Augmented Generation (RAG), which expands knowledge access at inference time, LmLm constrains factual memorization during pretraining by externalizing entity-level knowledge into a database. Since facts are not internalized in model weights, they can be instantly modified or removed through database operations alone (§ 4.3)—capabilities that are difficult to achieve with post-hoc retrieval. Current RAG systems still encode facts in parameters, so modifying specific knowledge requires model retraining or careful prompting to override internalized information (Xie et al., 2023; Hsia et al., 2024; Tan et al., 2024).
Model FactScore $\uparrow$ T-REx EM $\uparrow$ PopQA Acc $\uparrow$
OpenAI/GPT2-355M 15.2 28.4 19.1
   + RAG 20.1 75.8 37.5
GPT2-355M-LmLm 23.9+3.8 58.7-17.1 52.0+14.5
To contextualize their respective strengths, we compare against a controlled RAG baseline that retrieves the top-4 Wikipedia articles and prepends them to the standard LMs during generation. As shown in Table 3, LmLm achieves higher FactScore and PopQA accuracy at the current scale, indicating that it effectively queries its database for precise factual information. In contrast, RAG achieves stronger performance on T-REx, likely because the benchmark’s queries are drawn verbatim from Wikipedia, making retrieval from the same corpus nearly perfect.
These results point to complementary strengths: RAG enriches pretrained models (including LmLm) with broader contextual information at inference, whereas LmLm restructures pretraining itself to provide fine-grained factual grounding that is directly editable and verifiable. A hybrid system could retrieve documents with RAG for contextual understanding while using LmLm for precise, verifiable entity lookups within those documents.
5 Further Analysis
@@
Model Type FactScore ↑ T-REx EM ↑
@@
As shown in Figure 8, models trained with a standard SFT objective exhibit a rapid decrease in the loss on these tokens, suggesting memorization in the model’s parameters. In contrast, LmLm, trained with the masked loss maintains a high loss throughout training, indicating that these facts are not stored internally. LmLm’s successful application in the machine unlearning benchmark (Sec. 4.3) further supports this finding. Additionally, we observe a notable performance gap in Table 4, where factual precision drops substantially when the external database is disabled—forcing the model to rely on its internal parameters. These findings suggest that editing the database offers a direct way to control what the model knows and forgets.
How much knowledge should be externalized?
For our main experiments, we aim to exhaustively annotate all atomic entity-level knowledge. However, this may not be optimal. To explore the trade-off between storing knowledge in model parameters versus offloading it into the database, we devise a ranking criterion that quantifies the value of each lookup. Facts that require a lookup to achieve a high likelihood are difficult to learn—typically long-tail or highly specific knowledge—and should be externalized. On the other hand, externalizing facts that already have a high likelihood—typically common or easily inferred knowledge—provides limited upside and may harm general capabilities.
Concretely, we first train a LmLm model and a Standard model for one epoch on the data described in Sec. 4. For each fact, we then measure the average loss difference on its factual value tokens. A large difference indicates that Standard LMs struggle to learn the fact (See Appendix  B.3 for qualitative examples). Using this criterion, we pretrain LLMs with varying annotation ratios (e.g. For 90% knowledge offloading, we keep the top 90% knowledge with highest loss difference), interpolating between the two extremes of Standard (fully parametric) and LmLm (fully offloaded).
As shown in Figure 8, greater offloading consistently improves language modeling (lower perplexity) and factual precision, while preserving natural language understanding. This suggests that, for knowledge representable as triplets, increasing the offloading ratio is beneficial. Extending this spectrum to broader and more abstract forms of knowledge remains an open direction for future work.
Figure 7: Training loss on return value tokens.
Figure 8: Effect of different offloading ratios; within our scope, more offloading is beneficial.
6 Discussion
Toward Efficient Scaling via Knowledge Offloading.
We summarize our main findings in Figure 4. In particular, even smaller instances of LmLm match or exceed the performance of much larger off-the-shelf models. These results highlight the potential of LmLm to scale efficiently by offloading knowledge to an external database, thereby maintaining strong factual accuracy with fewer parameters. While our experiments are limited to modest scales due to computational constraints, the observed trends suggest that LmLms offer a promising direction toward parameter-efficient language models that externalize factual storage. Such models may enable real-time, verifiable knowledge updates and open up new possibilities for deploying LmLm in resource-constrained or fact-sensitive environments.
Beyond Entity-Level Knowledge Offloading.
By leveraging knowledge triplets, LmLm focuses on separating entity-level factual knowledge from linguistic competency. Current research on probing (Petroni et al., 2019; Hu et al., 2024), evaluating (Min et al., 2023; Wei et al., 2024), and editing (Meng et al., 2023a) internal knowledge in language models similarly concentrates on entity-level atomic facts. However, extending our method beyond entity-level knowledge remains a challenge. This difficulty arises in determining effective formats for separating knowledge beyond entity-focused triplets. For instance, using a simple QA function call format raises concerns about the potential for hallucinated facts during annotation.
@@ 6 Discussion
LmLm is a promising step towards separating factual knowledge from language models with many exciting future directions. Our current limitations include: (1) It does not guarantee perfect factuality during generation. Noise in the database and errors from fuzzy matching can introduce inaccuracies. However, such issues are easily traceable and verifiable for LmLm. (2) LmLm introduces additional tokens for lookup queries, which increases training and inference costs. (3) The current implementation focuses on entity-level factual knowledge, which captures only a subset of the broader knowledge spectrum. (4) Our experiments are limited to small models and datasets due to compute constraints. Although sufficient to show core benefits, scaling up may improve performance and support more complex tasks.
@@ Acknowledgments
This research was supported by a gift to the LinkedIn–Cornell Bowers Strategic Partnership. SZ is supported by the Defense Advanced Research Projects Agency (DARPA) under Grant No. D24AP00259-00. CKB is supported by the National Science Foundation (NSF) through the NSF Research Traineeship (NRT) program under Grant No. 2345579. JPZ is supported by a grant from the Natural Sciences and Engineering Research Council of Canada (NSERC) (567916). DG is supported by Empire AI Postdoctoral Fellowship. This work is partially supported by Open Philanthropy; the National Science Foundation NSF under awards OAC-2311521, IIS-2107161, IIS-1724282, and HDR-2118310; NASA under award No. 20-OSTFL20-0053; the Cornell Center for Materials Research with funding from the NSF MRSEC program (DMR-1719875); DARPA; arXiv; Apple; and the New York Presbyterian Hospital. We gratefully acknowledge use of the research computing resources of the Empire AI Consortium, Inc, with support from the State of New York, the Simons Foundation, and the Secunda Family Foundation (Bloom et al., 2025). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation or of NASA.
@@ A.1 Knowledge Extraction Details
We construct a high-quality seed dataset by sampling 1,000 passages from SQuAD-v2 (Rajpurkar et al., 2018) and 1,000 passages from Wikipedia. These passages are annotated by GPT-4o with structured factual triples and used to train the Annotator model. The remaining Wikipedia passages are used as the pre-training corpus for LmLm, with no overlap with the validation set.
@@
  • The Corrector is fine-tuned using LoRA (r=32, alpha=16(Hu et al., 2021) for 2 epochs on 19k GPT-4o-annotated SQuAD-v2 passages, with a learning rate of $2\times 10^{-4}$ and an effective batch size of 32. Sequence packing is enabled to improve training efficiency.
@@ A.3 Formalization of Training and Evaluation Objectives
  • $\mathcal{T}_{\text{train}}$: Return values and <|db_end|> are excluded from training loss objective.
@@
$\mathcal{T}_{\text{train}}$:
Napoleon was born on <|db_start|> Napoleon <|sep|> Birth_Date
<|db_retrieve|> August 15, 1769 <|db_end|> August 15, 1769.
@@
$\mathcal{L}(\theta)=-\sum_{t\in\mathcal{T}_{\text{train}}}\log p_{\theta}(x_{t}\mid x_{<t}),\quad\text{where }\mathcal{T}_{\text{train}}=\left\{t\mid x_{t}\notin\mathcal{T}_{\text{v}}\cup\{\texttt{<|db\_end|>}\}\right\}.$ (2)
@@
    $\text{PPL}_{\text{static/dynamic}}=\exp\left(-\frac{1}{|\mathcal{T}_{\text{org}}|}\sum_{t\in\mathcal{T}_{\text{org}}}\log p_{\theta}(x_{t}\mid x_{<t})\right).$
@@
    $\text{PPL}_{\text{norm}}=\exp\left(-\frac{1}{|\mathcal{T}_{\text{org}}|}\sum_{t\in\mathcal{T}_{\text{train}}}\log p_{\theta}(x_{t}\mid x_{<t})\right)$
@@
    $\text{NLL}(x)=-\sum_{t\in\mathcal{T}_{\text{train}}}\log p_{\theta}(x_{t}\mid x_{<t}).$
@@ A.4 Database and Retrieval Setting
We build our database by annotating the pre-training data, obtaining 54.6M knowledge triplets consisting of 9.5M unique entities, 8.5M relationships and 16.2M unique values. For retrieval, we employ a fuzzy matching mechanism based on the cosine similarity of sentence embeddings from all-MiniLM-L6-v2444https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. As shown in Figure 9, specifically, given a lookup call, we compute its embedding and compare it with the embeddings of stored triplets in our database. If the highest similarity score is below a threshold of 0.6, we return unknown to indicate that no sufficiently similar match was found. Alternatively, we implement a prefix-tree constrained generation, ensuring that lookup calls remain covered by the structured knowledge representations. See detailed discussion in Appendix D.3.